Please help! A little rusty with math...
A can of tomatoes is on sale at 20% off for 64 cents. What is the normal selling price of the tomatoes?
Now I know the answer is 80 cents, but I don't know how that answer was found.
When I try and solve, I come up with 76.8 cents. There has to be something I'm forgetting to do.
Thanks for your help.

Answer by Earlsdon(6294)

You can start by letting P = the normal selling price.
The sale price ($0.64) = the normal selling price (P) minus 20% (0.2) of the normal selling price.
So, P - 20% of P = $0.64 or, in algebrese:
P - 0.2P = $0.64 Simplify and solve for P.
0.8P = $0.64 Divide both sides by 0.8
P = $0.64/0.8 = $0.80 or 80 cents.